OpenStudy (anonymous):

Can someone please help with a few questions? For some of them I just need my answer checked. Thank you!! 1.The trunk of a pine tree will be cut up to make boards. •The trunk is 52 feet long. •The base of the trunk is a circle 1.5 feet across. •The trunk tapers gradually to a point at the top. Which expression best models the volume of the trunk? A. pi*.75^2*52 B. pi*1.5^2*52 C.1/3*pi*.75^2*52 D.1/3*pi*1.5^2*52 (my answer)

OpenStudy (anonymous):

Angle AEC is inscribed in circle D. Angle ABC circumscribes circle D. Which equation about the relationships in the circle is true? A. 1/2 m angle ADC = m Angle AEC B. 1/2 m angle ABC = m angle ADC C. 1/2 m angle ADC = m angle ABC D. 1/2 m angle ABC = m angle AEC

OpenStudy (anonymous):

3. The cone is completely filled with ice cream. There is a half sphere of ice cream on top. What is the total volume of ice cream in cubic inches? A. 6.8 B. 8.9 (my answer) C.35.6 D.52.4

OpenStudy (anonymous):

4. A waffle cone for ice cream has a diameter of 8 centimeters and a height of 12 centimeters. What is the volume of the cone, in cubic centimeters? A. 32π cubic centimeters B. 64π cubic centimeters C.192π cubic centimeters D.256π cubic centimeters 5. What is the volume of this square pyramid? A. 18 cubic inches B. 54 cubic inches C.108 cubic inches (my answer) D.324 cubic inches

OpenStudy (anonymous):

Segments CY and CZ are radii of circle C. Chord WZ is 5.6 centimeters long. What is the length of XY? A.0.6 centimeters B.1.4 centimeters C.2.1 centimeters D.2.8 centimeters

OpenStudy (anonymous):

What is the length of y on this circle? A. 8 units (my answer) B.10 units C.12 units D.18 units

OpenStudy (anonymous):

Which statement explains why all circles are similar? A.There are 360° in every circle. (my answer) B.The ratio of the circumference of a circle to its diameter is same for every circle. C.The diameter of every circle is proportional to the radius. D.The inscribed angle in every circle is proportional to the central angle. A square pyramid and a cone have the same base area. The volume of the cone is 10 cm3, and the height of the cone is 15 cm. The height of the square pyramid is 24 cm. Find the volume of the square pyramid to the nearest whole unit. A.144 cm3 B.160 cm3 C.480 cm3 D.2400 cm3